# Decomposing Global Feature Effects Based on Feature Interactions

### v2

@@ 10 Conclusion and Discussion
We introduced GADGET, a new framework that partitions the feature space into interpretable and distinct regions such that feature interactions of any subset of features are minimized. GADGET can be used with any feature effect method that satisfies the local decomposability axiom (Axiom 1). We demonstrated its use with well-known global feature effect methods—namely PD, ALE, and SD—and provided visualizations and estimates for regional feature effects and interaction-related heterogeneity. Furthermore, we introduced different measures to quantify and analyze feature interactions through the reduction of interaction-related heterogeneity within GADGET. To define the interacting feature subset, we introduced the PINT test, a novel permutation-based significance test to detect global feature interactions for any feature effect method used within GADGET.
@@
In general, our proposed method works well if learned feature interactions are not overly local and if observations can be grouped based on feature interactions such that local feature effects within the groups are homogeneous, while different groups exhibit heterogeneous feature effects. This approach helps to avoid the aggregation bias of global feature effect methods, provides deeper insights into learned regional effects, and detects potential biases within subgroups (as illustrated in Section 8). One of the real-world examples also showed that the frequently made assumption of symmetric feature interactions (e.g., in Shapley values) is not always the case (see also Masoomi et al., 2023, for research on asymmetrical feature interactions). Thus, including domain knowledge to define the interacting feature subset $Z$ might sometimes be meaningful.
Finally, GADGET is—as all other interpretation methods—affected by the Rashomon effect due to differently learned feature effects for different fitted models. Thus, the results of GADGET when applied to different models might not be the same, but might provide more insights when compared to each other. Analyzing the influence of the Rashomon effect is out-of-scope in this paper but will be considered in future work. More details and related work to this topic are provided in Appendix H.
@@
Appendix A Theoretical Evidence of GADGET
We provide the proofs for the theorems of Sections 4.1 to 4.5. This includes the theoretical foundation of GADGET as well as the proofs of the applicability of the feature effect methods PD, ALE, and SD within the GADGET algorithm by defining respective local feature effect functions that fulfill Axiom 1. We also show in Appendix A.4 that the REPID method which was introduced in Section 3 is a special case of the GADGET algorithm.
A.1 Proof of Theorem 2
@@
$\mathcal{L}\left(\mathcal{A}_{g},x_{j}\right)$ $=$ $\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{g}}\left(h(x_{j},\mathbf{x}_{-j}^{(i)})-\mathds{E}[h(x_{j},X_{-j})|\mathcal{A}_{g}]\right)^{2}$
@@
A.2 Proof of Theorem 3
Proof Sketch We show that the theoretical minimum of the objective is $\mathcal{I}(t^{*},z^{*})=0$ if no feature in $S$ interacts with any feature in $-Z$. We apply the functional decomposition within the risk function $\mathcal{R}(\mathcal{A}_{b}^{t,z},\tilde{\mathbf{x}}_{j})$. Since we assume that no feature in $S$ interacts with any feature in $-Z$, the second and third term—which contain interactions between these two feature subsets—are zero. Hence, the risk function is only defined by feature interactions between features in $S$ and $Z$, which are minimized by the objective in Algorithm 1.
@@ A.2 Proof of Theorem 3
$\mathcal{R}(\mathcal{A}_{b}^{t,z},\tilde{\mathbf{x}}_{j})$ $=$ $\sum\limits_{\begin{subarray}{c}k:k\in\{1,\ldots,m\}\\ \wedge x_{j}^{(k)}\in\mathcal{A}_{b}^{t,z}\end{subarray}}\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{b}^{t,z}}\left(\sum\limits_{l=1}^{p-1}\;\sum\limits_{\begin{subarray}{c}W\subseteq-j,\\ |W|=l\end{subarray}}g_{W\cup j}(x_{j},\mathbf{x}_{W}^{(i)})-\mathds{E}[g_{W\cup j}(x_{j},X_{W})|\mathcal{A}_{b}^{t,z}]\right)^{2}$
@@
$\mathcal{R}(\mathcal{A}_{b}^{t,z},\tilde{\mathbf{x}}_{j})$ $=$ $\sum\limits_{\begin{subarray}{c}k:k\in\{1,\ldots,m\}\\ \wedge x_{j}^{(k)}\in\mathcal{A}_{b}^{t,z}\end{subarray}}\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{b}^{t,z}}\left(\Bigg{(}\sum\limits_{l=1}^{|Z\setminus j|}\;\sum\limits_{\begin{subarray}{c}Z_{l}\subseteq Z\setminus j,\\ |Z_{l}|=l\end{subarray}}g_{Z_{l}\cup j}(x_{j},\mathbf{x}_{Z_{l}}^{(i)})-\mathds{E}[g_{Z_{l}\cup j}(x_{j},X_{Z_{l}})|\mathcal{A}_{b}^{t,z}]\Bigg{)}\right.$
@@ A.3 Proof of Theorem 4
A.4 REPID as Special Case of GADGET
The objective function $\mathcal{I}(t,z)$ in Algorithm 1 for $h=\hat{f}^{c}(x_{j},\mathbf{x}_{-j}^{(i)})$ is defined by the above loss function $\mathcal{L}^{PD}\left(\mathcal{A}_{g},x_{j}\right)$ as follows:
$\mathcal{I}(t,z)=\sum_{j\in S}\sum_{g\in\{l,r\}}\sum_{k:k\in\{1,\ldots,m\}\wedge\mathbf{x}_{j}^{(k)}\in\mathcal{A}_{g}}\mathcal{L}^{PD}\left(\mathcal{A}_{g},\mathbf{x}_{j}^{(k)}\right)$
@@ A.4 REPID as Special Case of GADGET
$\mathcal{I}(t,z)=\sum_{g\in\{l,r\}}\sum_{k=1}^{m}\mathcal{L}^{PD}\left(\mathcal{A}_{g},\mathbf{x}_{j}^{(k)}\right),$
@@ A.6 Proof of Theorem 6
Appendix B Further Characteristics of Feature Effect Methods
@@ B.3 Decomposability of SD
B.4 Overview on Estimates and Visualizations
@@
  • PD: $95\%$ interval for (mean-centered) regional PD $\Big{[}\hat{f}^{PD,c}_{j|\mathcal{A}_{g}}(x_{j})\pm 1.96\cdot\sqrt{\hat{\mathcal{L}}^{PD}(\mathcal{A}_{g},x_{j})}\Big{]}$.
  • ALE: standard deviation of local effects ^$\sqrt{\hat{\mathcal{L}}^{ALE}(\mathcal{A}_{g},x_{j})}$.
@@
Note that we can also visualize the non-centered PD ($\hat{f}^{PD}_{j|\mathcal{A}_{g}}$) instead of the mean-centered PD, which might provide more insights regarding interpretation. However, the interaction-related heterogeneity must be estimated by the mean-centered ICE curves to only represent heteroegeneity induced by feature interactions (see Appendix A.3).
B.5 Handling of Categorical Features
@@
Appendix C Definition and Specification of Stop Criteria for GADGET
@@
C.1 Definition of Stop Criteria
We propose the following stopping criteria to control the number of partitioning steps in GADGET: First, we can use common decision tree hyperparameters such as the tree depth or the minimum number of observations per leaf node. Alternatively, an early stop mechanism can be applied based on the interaction-related heterogeneity reduction. According to our proposed split-wise measure, we further split only if the relative improvement is at least $\gamma\in[0,1]$ times the total relative heterogeneity reduction of the previous split: $\gamma\times\frac{\sum_{j\in S}\left(\mathcal{R}(\mathcal{A}_{P},\tilde{\mathbf{x}}_{j})\right)-\mathcal{I}(\hat{t},\hat{z})}{\sum_{j\in S}\left(\mathcal{R}(\mathcal{X},\tilde{\mathbf{x}}_{j})\right)}$. Another option is to stop splitting once a pre-defined total reduction of heterogeneity ($R^{2}_{Tot}$) is achieved. Generally, higher $\gamma$ values and lower thresholds for $R^{2}_{Tot}$ result in fewer partitioning steps, and vice versa.
C.2 Recommendations to Specify Stop Criteria
GADGET performs binary splits, making it easier to minimize abrupt interactions (i.e., interactions where categorical or discretized features are involved, see also Section B.2). Conversely, minimizing smooth interactions (i.e., interactions between continuous features) may require multiple splits. We recommend a maximum tree depth between 3 and 6 to maintain some degree of interpretability and setting $\gamma$ (the minimum relative risk reduction for a split) between 0.1 and 0.25 to further limit the number of terminal regions. For abrupt interactions (with few interactions), the risk reduction for the first split is usually quite high. Therefore, to find further feature interactions, we recommend to set $\gamma$ towards the lower end of the recommended range (e.g., 0.1 to 0.2). In the case of smooth interactions, the risk reduction for the first split is usually rather low compared to subgroup-specific feature interactions. Therefore, setting $\gamma$ to a small value leads to a vanishing improvement which prevents the desired stop mechanism. However, the value should also not be too high to prevent the first split. In summary, a tree depth of 3 to 6 and $\gamma$ between 0.1 and 0.25 worked well in our examples. To precisely quantify feature interactions or to completely minimize them, more splits and a smaller $\gamma$ may be needed, possibly using $\mathcal{R}^{2}_{Tot}$ as a stopping criterion.
Appendix D Theoretical Evidence of PINT
@@
D.1 Proof of Theorem 7
Proof Sketch In this proof we show that the type I error $\mathds{P}(\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha})$ of the test is constrained by $\alpha+\epsilon$ with $\epsilon=\max_{t>0}[\mathds{P}(\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})>t)-\mathds{P}(\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>t)]$.
@@ D.1 Proof of Theorem 7
$\quad\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha}\biggr{)}$ (37)
$=\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha}\biggr{)}+\biggl{[}\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha}\biggr{)}-\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha}\biggr{)}\biggr{]}$ (38)
$\leq\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha}\biggr{)}+\max_{t>0}\biggl{[}\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})>t\biggr{)}-\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>t\biggr{)}\biggr{]}$ (39)
@@
D.2 Proof of Theorem 8
Proof Sketch Here, we show that the type II error of the PINT test strives towards 00. Recall that a risk of 0 means that no interactions are learned by the model. If Condition (18) holds—i.e., the risk under the permuted data set is less than every $\epsilon>0$ with high probability—the type II error of the PINT test strives towards 00. The last step follows from the assumption that feature interactions have been learned by the model fitted on the original data set. It follows that the power of the test strives towards $1$.

### v3

@@ 10 Conclusion and Discussion
We introduced GADGET, a new framework that partitions the feature space into interpretable and distinct regions such that feature interactions of any subset of features are minimized. GADGET can be used with any feature effect method that satisfies the local decomposability axiom (Axiom 1). We demonstrated its use with well-known global feature effect methods—namely PD, ALE, and SD—and provided visualizations and estimates for regional feature effects and interaction-related heterogeneity. Furthermore, we introduced different measures to quantify and analyze feature interactions through the reduction of interaction-related heterogeneity within GADGET. To define the interacting feature subset, we introduced the novel PINT procedure to detect global feature interactions for any feature effect method used within GADGET.
@@
In general, our proposed method works well if learned feature interactions are not overly local and if observations can be grouped based on feature interactions such that local feature effects within the groups are homogeneous, while different groups exhibit heterogeneous feature effects. This approach helps to avoid the aggregation bias of global feature effect methods, provides deeper insights into learned regional effects, and detects potential biases within subgroups (as illustrated in Section 8). One of the real-world examples also showed that the frequently made assumption of symmetric feature interactions (e.g., in Shapley values) is not always the case (see also Masoomi et al., 2022, for research on asymmetrical feature interactions). Thus, including domain knowledge to define the interacting feature subset $Z$ might sometimes be meaningful.
Finally, GADGET is—as all other interpretation methods—affected by the Rashomon effect due to differently learned feature effects for different fitted models. Thus, the results of GADGET when applied to different models might not be the same, but might provide more insights when compared to each other. Analyzing the influence of the Rashomon effect is out-of-scope in this paper but will be considered in future work. More details and related work to this topic are provided in Appendix I.
@@
Appendix A Further Details on Functional ANOVA Decomposition
As stated in Section 2.2, the functional ANOVA provides a decomposition of the prediction function into the main and higher-order effect of all involved features (Sobol’, 1993; Stone, 1994; Hooker, 2004, 2007; Li and Rabitz, 2012):
$\hat{f}(\mathbf{x})=g_{0}+\sum_{j=1}^{p}g_{j}(\mathbf{x}_{j})+\sum_{j\neq k}g_{jk}(\mathbf{x}_{j},\mathbf{x}_{k})+\ldots+g_{12\ldots p}(\mathbf{x})=g_{0}+\sum_{k=1}^{p}\;\sum_{\begin{subarray}{c}W\subseteq\{1,\ldots,p\},\\ |W|=k\end{subarray}}g_{W}(\mathbf{x}_{W}),$
where $g_{0}$ is a constant, $g_{j}(\mathbf{x}_{j})$ denotes the main effect of the $j$-th feature, $g_{jk}(\mathbf{x}_{j},\mathbf{x}_{k})$ is the pure two-way interaction effect between features $\mathbf{x}_{j}$ and $\mathbf{x}_{k}$, and so forth. The last component $g_{12\ldots p}(\mathbf{x})$ always allows for an exact decomposition. We discuss here the standard and the generalized functional ANOVA decomposition which provide a unique decomposition of the above-stated equation based on different assumptions.
Standard Functional ANOVA Decomposition.
[…]
Generalized Functional ANOVA Decomposition.
[…]
With $p_{X}(\mathbf{x})$ now being a general probability density with its support being grid-closed and with $\hat{f}$ being square-integrable, the component functions can then be uniquely determined by optimizing Eq. (21).
[…]
The estimation procedure can become computationally expensive with increasing dimensionality. However, being able to decompose a prediction function into its components and hence, being able to interpret main and higher-order effects between features separately is often desired. Thus, models that are able to directly model these structures have been introduced and extended. Generalized additive models (GAMs), for example, additively decompose the prediction function into the (non-linear) main effects of each feature. Lou et al. (2013) extended GAMs by adding relevant two-way feature interactions to the main effect model. Watanabe et al. (2021) further extend this approach by also allowing feature interactions of a higher order as well as monotonic constraints.
Recent research has focused on computing the decomposition more efficiently, e.g., by developing models that integrate the respective conditions directly in the optimization process through constraints (e.g., Sun et al., 2022) or on purifying the resulting decomposition for specific models within the modeling process (e.g., Lengerich et al., 2020; Hu et al., 2022). However, most approaches require estimating the underlying data distribution which remains challenging (Lengerich et al., 2020), with initial attempts employing adaptive kernel methods (Sun et al., 2022). Notably, finding an appropriate decomposition is only complex in the presence of feature interactions, otherwise, the prediction function can be uniquely decomposed into the main effects using models like the generalized additive model (GAM).
Appendix B Theoretical Evidence of GADGET
We provide the proofs for the theorems of Sections 4.1 to 4.5. This includes the theoretical foundation of GADGET as well as the proofs of the applicability of the feature effect methods PD, ALE, and SD within the GADGET algorithm by defining respective local feature effect functions that fulfill Axiom 1. We also show in Appendix B.4 that the REPID method which was introduced in Section 3 is a special case of the GADGET algorithm.
B.1 Proof of Theorem 2
@@
$\mathcal{L}_{j}\left(\mathcal{A}_{g},x_{j}\right)$ $=$ $\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{g}}\left(h(x_{j},\mathbf{x}_{-j}^{(i)})-\mathds{E}[h(x_{j},X_{-j})|\mathcal{A}_{g}]\right)^{2}$
@@
B.2 Proof of Theorem 3
Proof Sketch We show that the theoretical minimum of the objective is $\mathcal{I}(t^{*},z^{*})=0$ if no feature in $S$ interacts with any feature in $-Z$. We apply the functional decomposition within the risk function $\mathcal{R}_{j}(\mathcal{A}_{b}^{t,z},\tilde{\mathbf{x}}_{j})$. Since we assume that no feature in $S$ interacts with any feature in $-Z$, the second and third term—which contain interactions between these two feature subsets—are zero. Hence, the risk function is only defined by feature interactions between features in $S$ and $Z$, which are minimized by the objective in Algorithm 1.
@@ B.2 Proof of Theorem 3
$\mathcal{R}_{j}(\mathcal{A}_{b}^{t,z},\tilde{\mathbf{x}}_{j})$ $=$ $\sum\limits_{\begin{subarray}{c}k:k\in\{1,\ldots,m\}\\ \wedge x_{j}^{(k)}\in\mathcal{A}_{b}^{t,z}\end{subarray}}\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{b}^{t,z}}\left(\sum\limits_{l=1}^{p-1}\;\sum\limits_{\begin{subarray}{c}W\subseteq-j,\\ |W|=l\end{subarray}}g_{W\cup j}(x_{j},\mathbf{x}_{W}^{(i)})-\mathds{E}[g_{W\cup j}(x_{j},X_{W})|\mathcal{A}_{b}^{t,z}]\right)^{2}$
@@
$\mathcal{R}_{j}(\mathcal{A}_{b}^{t,z},\tilde{\mathbf{x}}_{j})$ $=$ $\sum\limits_{\begin{subarray}{c}k:k\in\{1,\ldots,m\}\\ \wedge x_{j}^{(k)}\in\mathcal{A}_{b}^{t,z}\end{subarray}}\sum\limits_{i:\mathbf{x}^{(i)}\in\mathcal{A}_{b}^{t,z}}\left(\Bigg{(}\sum\limits_{l=1}^{|Z\setminus j|}\;\sum\limits_{\begin{subarray}{c}Z_{l}\subseteq Z\setminus j,\\ |Z_{l}|=l\end{subarray}}g_{Z_{l}\cup j}(x_{j},\mathbf{x}_{Z_{l}}^{(i)})-\mathds{E}[g_{Z_{l}\cup j}(x_{j},X_{Z_{l}})|\mathcal{A}_{b}^{t,z}]\Bigg{)}\right.$
@@ B.3 Proof of Theorem 4
B.4 REPID as Special Case of GADGET
The objective function $\mathcal{I}(t,z)$ in Algorithm 1 for $h=\hat{f}^{c}(x_{j},\mathbf{x}_{-j}^{(i)})$ is defined by the above loss function $\mathcal{L}_{j}^{PD}\left(\mathcal{A}_{g},x_{j}\right)$ as follows:
$\mathcal{I}(t,z)=\sum_{j\in S}\sum_{g\in\{l,r\}}\sum_{k:k\in\{1,\ldots,m\}\wedge\mathbf{x}_{j}^{(k)}\in\mathcal{A}_{g}}\mathcal{L}_{j}^{PD}\left(\mathcal{A}_{g},\mathbf{x}_{j}^{(k)}\right)$
@@ B.4 REPID as Special Case of GADGET
$\mathcal{I}(t,z)=\sum_{g\in\{l,r\}}\sum_{k=1}^{m}\mathcal{L}_{j}^{PD}\left(\mathcal{A}_{g},\mathbf{x}_{j}^{(k)}\right),$
@@ B.6 Proof of Theorem 6
Appendix C Further Characteristics of Feature Effect Methods
@@ C.3 Decomposability of SD
C.4 Overview on Estimates and Visualizations
@@
  • PD: $95\%$ interval for (mean-centered) regional PD $\Big{[}\hat{f}^{PD,c}_{j|\mathcal{A}_{g}}(x_{j})\pm 1.96\cdot\sqrt{\hat{\mathcal{L}_{j}}^{PD}(\mathcal{A}_{g},x_{j})}\Big{]}$.
  • ALE: standard deviation of local effects $\sqrt{\hat{\mathcal{L}_{j}}^{ALE}(\mathcal{A}_{g},x_{j})}$.
@@
Note that we can also visualize the non-centered PD ($\hat{f}^{PD}_{j|\mathcal{A}_{g}}$) instead of the mean-centered PD, which might provide more insights regarding interpretation. However, the interaction-related heterogeneity must be estimated by the mean-centered ICE curves to only represent heteroegeneity induced by feature interactions (see Appendix B.3).
C.5 Handling of Categorical Features
@@
Appendix D Definition and Specification of Stop Criteria for GADGET
@@
D.1 Definition of Stop Criteria
We propose the following stopping criteria to control the number of partitioning steps in GADGET: First, we can use common decision tree hyperparameters such as the tree depth or the minimum number of observations per leaf node. Alternatively, an early stop mechanism can be applied based on the interaction-related heterogeneity reduction. According to our proposed split-wise measure, we further split only if the relative improvement is at least $\gamma\in[0,1]$ times the total relative heterogeneity reduction of the previous split: $\gamma\times\frac{\sum_{j\in S}\left(\mathcal{R}_{j}(\mathcal{A}_{P},\tilde{\mathbf{x}}_{j})\right)-\mathcal{I}(\hat{t},\hat{z})}{\sum_{j\in S}\left(\mathcal{R}_{j}(\mathcal{X},\tilde{\mathbf{x}}_{j})\right)}$. Another option is to stop splitting once a pre-defined total reduction of heterogeneity ($R^{2}_{Tot}$) is achieved. Generally, higher $\gamma$ values and lower thresholds for $R^{2}_{Tot}$ result in fewer partitioning steps, and vice versa.
D.2 Recommendations to Specify Stop Criteria
GADGET performs binary splits, making it easier to minimize abrupt interactions (i.e., interactions where categorical or discretized features are involved, see also Section C.2). Conversely, minimizing smooth interactions (i.e., interactions between continuous features) may require multiple splits. We recommend a maximum tree depth between 3 and 6 to maintain some degree of interpretability and setting $\gamma$ (the minimum relative risk reduction for a split) between 0.1 and 0.25 to further limit the number of terminal regions. For abrupt interactions (with few interactions), the risk reduction for the first split is usually quite high. Therefore, to find further feature interactions, we recommend to set $\gamma$ towards the lower end of the recommended range (e.g., 0.1 to 0.2). In the case of smooth interactions, the risk reduction for the first split is usually rather low compared to subgroup-specific feature interactions. Therefore, setting $\gamma$ to a small value leads to a vanishing improvement which prevents the desired stop mechanism. However, the value should also not be too high to prevent the first split. In summary, a tree depth of 3 to 6 and $\gamma$ between 0.1 and 0.25 worked well in our examples. To precisely quantify feature interactions or to completely minimize them, more splits and a smaller $\gamma$ may be needed, possibly using $R^{2}_{Tot}$ as a stopping criterion.
Appendix E Theoretical Evidence of PINT
@@
E.1 Proof of Theorem 7
Proof Sketch In this proof we show that the type I error $\mathds{P}(\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha})$ of the test is constrained by $\alpha+\epsilon$ with $\epsilon=\max_{t>0}[\mathds{P}(\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})>t)-\mathds{P}(\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>t)]$.
@@ E.1 Proof of Theorem 7
$\quad\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha}\biggr{)}$ (39)
$=\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha}\biggr{)}+\biggl{[}\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha}\biggr{)}-\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha}\biggr{)}\biggr{]}$ (40)
$\leq\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>z_{1-\alpha}\biggr{)}+\max_{t>0}\biggl{[}\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})>t\biggr{)}-\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>t\biggr{)}\biggr{]}$ (41)
@@
E.2 Proof of Theorem 8
Proof Sketch Here, we show that the type II error of the PINT procedure strives towards 00. Recall that a risk of 0 means that no interactions are learned by the model. If Condition (18) holds—i.e., the risk under the permuted data set is less than every $\epsilon>0$ with high probability—the type II error of the PINT procedure strives towards 00. The last step follows from the assumption that feature interactions have been learned by the model fitted on the original data set. It follows that the power of the test strives towards $1$.